OpenStudy (anonymous):

If 2^(16x^2-8x-3)=1, what does x equal?

OpenStudy (anonymous):

Explanation please

OpenStudy (anonymous):

can somebody please help me with my problem? sorry to bother

OpenStudy (anonymous):

\[\large 2^{(16x^2-8x-3)}=1\]\[\large \log_22^{(16x^2-8x-3)}=\log_21\]\[\large (16x^2-8x-3) \log_22=\log_21\]\[\large \log_22=1, \space \log_21=0\]\[\large 16x^2-8x-3=0\]Now factor to solve for x or use the quadratic (suggested method)

OpenStudy (anonymous):

-1/4 and 3/4?

OpenStudy (anonymous):

@ChmE Am i right?

OpenStudy (anonymous):

I don't know I haven't did it

OpenStudy (anonymous):

Seems right !!

OpenStudy (anonymous):

cool,thanks for the help :) Happy New Years!

OpenStudy (anonymous):

Post the quick way then plz @Hero

OpenStudy (anonymous):

That is true, at least now they see how to do it, if that wasn't recognized right away